# The Complexity of Proper Homotopy Equivalence of Graphs

### v1

@@ 2.3. Surfaces
By surface we mean a second countable, Hausdorff, orientable, 2-manifold without boundary. A surface is of finite type if its fundamental group is finitely generated. Two finite-type surfaces are homeomorphic if they have the same number of genus and punctures. A surface is of infinite type if its fundamental group is not finitely generated. The classification of infinite-type surfaces predates, but is in direct analogy with, the classification of infinite graphs up to proper homotopy equivalence. It is due independently to Kerékjártó [Ker23] and Richards [Ric63].
@@
In this statement, $g(\Sigma)\in\mathbb{N}\cup\{0,\infty\}$ is the number of genus, $E(\Sigma)$ is the end space, and $E_{g}(\Sigma)$ is the subspace of ends accumulated by genus. The endspace of a surface can be defined mutatis mutandis with 2.1, replacing ${\Gamma}$ with $\Sigma$. To get the ends accumulated by genus, replace rank with genus in the definition of ends accumulated by loops.
2.3.1. Spaces of Surfaces
Next we want to realize sets of noncompact surfaces as Polish spaces; this includes infinite-type surfaces as well as finite-type surfaces with punctures. First, we define a pair of pants to be any orientable genus zero surface with at least one boundary component. If the pair of pants $P$ has $d+1$ boundary components we call them specifically a $d$-legged pair of pants, so that a $3$-holed sphere has two legs (the third boundary component being the waist). We allow $1$-legged pants, which are topologically annuli, and 00-legged pants, which are topologically disks. A pants decomposition of a surface $S$ is a collection of subsurfaces $\mathcal{P}=\{P_{j}\}_{j\in I}$ which are each themselves pants, cover the surface $S$, and overlap exactly on their boundary components. Denote the boundary components of $P_{j}$ by $\{\partial_{1}P_{j},\dots,\partial_{d+1}P_{j}\}$. That is, $S$ can be constructed by gluing the pants together along their boundary components by orientation preserving homeomorphisms, and we can write
@@ 2.3.1. Spaces of Surfaces
For any surface with a pants decomposition $(S,\mathcal{P})$ we associate a graph in the following way. For each pair of paints $P_{j}$, associate a vertex $v_{j}$. Whenever a boundary component of $P_{j}$ is glued to a boundary component of $P_{k}$, associate an edge from $(v_{j},v_{k})$, note that $j$ and $k$ may be equal. In this way, each $d$-legged pants corresponds to a vertex of degree $d+1$. If we consider surfaces with a basepoint interior to the pants decomposition, then we can associate a based graph. See fig. 6 for an example.
@@
For each space of graphs $\mathcal{G}_{\star}$ discussed in Section 2.2 we will define an analogous space of surfaces, $\mathcal{S}_{\star}$. The maps $\Phi_{\star}:\hat{\mathcal{S}_{\star}}\rightarrow\mathcal{G}_{\star}$ are surjective, so we can then endow $\hat{\mathcal{S}_{\star}}$ with the initial topology. Recall that the initial topology is the coarsest topology which makes $\Phi_{\star}$ continuous and has the sets
=$\mathcal{B}=\{\Phi^{-1}(\mathcal{U})\mid\mathcal{U}\text{ is open in }\mathcal{G}_{\star}\}$
@@
as a subbase. Note that this topology will not be Hausdorff. For example, the same surface with two pants decompositions in which pants curves are isotopic, but not equal, are indistinguishable by open sets in the topology. So, we quotient by the equivalence relation induced by $\Phi_{\star}$, that is $(S,\mathcal{P},b)\sim(S^{\prime},\mathcal{P}^{\prime},b^{\prime})$ if $\Phi_{\star}(S,\mathcal{P},b)=\Phi_{\star}(S^{\prime},\mathcal{P}^{\prime},b^{\prime})$. Call these quotient spaces (with the quotient topologies) $\mathcal{S}_{\star}$, and note that they are Hausdorff and homeomorphic to the corresponding space $\mathcal{G}_{\star}$. For detailed explanation, see for example Corollary 22.3 of Munkres’ Topology, [Mun00].
@@
Points in $\mathcal{S}_{\star}$ can be thought of as surfaces, up to homeomorphism, with based pants decompositions, up to isotopy. A puncture on a surface will map to a ray in the associated graph. So, similar to the discussion of which graphs live in which spaces $\mathcal{G}_{\star}$, there are no surfaces with isolated punctures in $\mathcal{S}_{k}$ and all homeomorphism types of noncompact surfaces are realized in $\mathcal{S}_{\leq k}$ for $k\geq 3$.
2.4. Complexity of Borel equivalence relations
Our main result, 3.2, discusses the complexity of the proper homotopy equivalence relation on $\mathcal{G}_{n}$ for $n\geq 3$. To formalize what we mean by complexity, we need to introduce the notion of Borel reducibility of equivalence relations. For a more in depth introduction to this area, we refer the reader to [Tse22, Part 4].
By a Borel (resp., analytic) equivalence relation $E$ on a Polish space $X$, we mean that $E$ is a Borel (resp., analytic) subset of $X^{2}$ with respect to the product topology. For equivalence relations $E$ and $F$ on Polish spaces $X$ and $Y$, we say that $E$ is Borel reducible to $F$, and we write $E\leq_{B}F$, if there is a Borel function $\phi:X\to Y$ so that $(x,y)\in E$ if and only if $(\phi(x),\phi(y))\in F$. In other words, to check whether two points are $E$-equivalent, we can pass the points to $Y$ via our Borel function and check whether the images are $F$-equivalent. In this case, we say that $\phi$ is a Borel reduction from $E$ to $F$. We say that $E$ is Borel bireducible with $F$ (or that $E$ and $F$ are Borel equivalent) if $E\leq_{B}F$ and $F\leq_{B}E$.
An equivalence relation $E$ is called concretely classifiable (or smooth) if there is a Borel reduction from $E$ to the equality relation on some Polish space $Y$. For example, similarity of $n\times n$ matrices is smooth since similarity is determined by computing the Jordan canonical form. However, many notable equivalence relations are not smooth. We say that $E$ is classifiable by countable structures if for some countable language $\mathcal{L}$, there is a Borel reduction from $E$ to the isomorphism relation of the space of $\mathcal{L}$-structures with universe $\mathbb{N}$.
By [BK96], there is an orbit equivalence relation, denoted $C_{\infty}$, induced by a Borel action of $S_{\infty}$ (the Polish group of all permutations of $\mathbb{N}$) that is “complete” in the sense that any orbit equivalence relation induced by a Borel action of a closed subgroup of $S_{\infty}$ is Borel reducible to $C_{\infty}$. In the literature, analytic equivalence relations which are bireducible with $C_{\infty}$ are often referred to as Borel complete. We will see in 3.2 that the proper homotopy equivalence relation on each of the spaces $\mathcal{G}_{\star}$ is Borel complete.
An important class of analytic equivalence relations that are reducible to $C_{\infty}$ are those that are classifiable by countable structures: if $\mathcal{L}$ is a countable language, then the isomorphism relation of the space of $\mathcal{L}$-structures can be viewed as an $S_{\infty}$ orbit equivalence relation (see, e.g. [Kec95, 16.C]).
3. Proofs of main results for 3-regular graphs
We begin by showing that every closed subset of the Cantor set can be found as the endspace of a graph in $\mathcal{G}_{3}$. This allows us to define a Borel reduction from the homeomorphism relation on closed subsets of the Cantor set to the PHE relation.
Lemma 3.1.
For any closed subset $C$ of the Cantor space, $C$ is realized as the end space accumulated by loops of some $\Gamma_{C}\in\mathcal{G}_{3}$. Moreover, $\Gamma_{C}$ can be taken to have all of its ends accumulated by loops.
@@ 3. Proofs of main results for 3-regular graphs
We use the standard identification of $2^{\mathbb{N}}$ with the Cantor set and with ends of a rooted binary tree. Let $C\subseteq{2^{\mathbb{N}}}$ be closed and not a singleton. Let ${\Gamma}_{C}^{*}$ be the subgraph of the rooted binary tree with endspace $C$ and no leaves. Note that every vertex in $\Gamma_{C}^{*}$ has valence $2$ or $3$. Let $\Gamma_{C}$ be the graph obtained by subdividing each edge of ${\Gamma}_{C}^{*}$ and then adding “lollipops” to each vertex of degree two, as illustrated in fig. 7. This modification doesn’t add or collapse any ends, so the resulting graph $\Gamma_{C}$ has $E({\Gamma_{C}})=E_{\ell}({\Gamma_{C}})=C$, and is of infinite rank.
Figure 7. Modifying a graph $\Gamma_{C}^{*}$ (left) so that it is in $\mathcal{G}_{3}$ by subdividing edges (middle) and then adding lollipops to degree $2$ vertices (right).
If $C$ is a singleton, then an example of a graph ${\Gamma}_{C}$ in $\mathcal{G}_{3}$ is given in fig. 8.
We are now ready to prove our main theorem for the space $\mathcal{G}_{3}$.
Theorem 3.2.
Proper homotopy equivalence on $\mathcal{G}_{3}$ is bireducible with $C_{\infty}$. That is, $\operatorname{\thicksim_{\text{PHE}}}$ is Borel complete.
@@
We start by showing that $C_{\infty}\leq_{B}\;\operatorname{\thicksim_{\text{PHE}}}$. By [CG01], it suffices to demonstrate a Borel reduction from the homeomorphism relation for closed subsets of the Cantor space to $\operatorname{\thicksim_{\text{PHE}}}$ on $\mathcal{G}_{3}$. This follows from 3.1 and the observation that the assignment $C\mapsto\Gamma_{C}$ in the proof of 3.1 is Borel. Indeed, because each nonempty basic open subset of $\mathcal{G}_{3}$ is determined by finitely many conditions on the graph; hence, the preimage of the set is Borel in $\mathcal{K}(2^{\mathbb{N}})$. Hence, the map $\phi:\mathcal{K}(2^{\mathbb{N}})\to\mathcal{G}_{3}$ defined by $C\mapsto\Gamma_{C}$ is a Borel reduction of homeomorphisms of closed subsets of $2^{\mathbb{N}}$ to $\operatorname{\thicksim_{\text{PHE}}}$, as desired.
Next we show the other direction, that $\operatorname{\thicksim_{\text{PHE}}}\;\leq_{B}\;C_{\infty}$; we give the intuition before the formal argument.
Intuitively: 2.2 tells us that two graphs are proper homotopy equivalent when they have the same rank and simultaneously homeomorphic endspace pairs. Each part of the endspace pair is homeomorphic to a closed subset of $2^{\mathbb{N}}$, the Cantor set. By [CG01], the homeomorphism relation on closed subsets of $2^{\mathbb{N}}$ is bireducible with isomorphism of countable structures. Since homeomorphism is the most complicated aspect of $\operatorname{\thicksim_{\text{PHE}}}$, we have the desired Borel reduction.
Formally: We use a similar argument to the one in [CG01]. Let $\mathcal{K}(2^{\mathbb{N}})$ be the space of closed subsets of the Cantor space, equipped with the Vietoris topology. Let $\mathcal{L}$ be the language of Boolean algebras $\{\vee,\wedge,-,0,1\}$, along with an extra constant symbol $n$, unary relation symbols $K$ and $L$, and unary function symbol $f$. Let $\mathcal{T}$ be the axioms of Boolean algebras along with a sentence asserting that $K$ and $L$ are disjoint, and that $f$ is a surjection from $K$ to $L$. Let $\mathcal{M}$ be the space of models of $\mathcal{T}$ with universe $\mathbb{N}$.
To each $\Gamma\in\mathcal{G}_{k}$ with rank $\operatorname{rk}(\Gamma)\in\mathbb{N}\cup\{\mathbb{N}\}$, end space $E(\Gamma)$, and end space accumulated by loops $E_{\ell}(\Gamma)$, we associate the countable structure $\mathcal{B}_{\Gamma}\in\mathcal{M}$:
$(n_{\Gamma},K_{\Gamma},L_{\Gamma},f_{\Gamma},\vee,\wedge,-,0,1)$
where:
  • $n_{\Gamma}=\text{rk}_{\Gamma}+1$ if $\text{rk}_{\Gamma}\in\mathbb{N}$ or $n_{\Gamma}=0$ if $\text{rk}_{\Gamma}=\mathbb{N}$,
  • $K_{\Gamma}$ is the Boolean algebra of clopen subsets of $E(\Gamma)$,
  • $L_{\Gamma}$ is the Boolean algebra of clopen subsets of $E_{\ell}(\Gamma)$,
  • $f_{\Gamma}$ is the (surjective) map $f_{\Gamma}:K_{\Gamma}\to L_{\Gamma}$ defined by $Y\mapsto Y\cap E_{\ell}(\Gamma)$,
  • and $\vee,\wedge,-,0,1$ are the usual Boolean operations.
By the usual proof of Stone duality, $\Gamma_{1}\operatorname{\thicksim_{\text{PHE}}}\Gamma_{2}$ if and only if $\mathcal{B}_{\Gamma_{1}}\cong\mathcal{B}_{\Gamma_{2}}$. To see that the assignment $\Gamma\mapsto\mathcal{B}_{\Gamma}$ is Borel, we first note that the assignments $\Gamma\mapsto n_{\Gamma}\in\mathbb{N}$, $\Gamma\mapsto E(\Gamma)\in\mathcal{K}(2^{\mathbb{N}})$, and $\Gamma\mapsto E_{\ell}(\Gamma)\in\mathcal{K}(2^{\mathbb{N}})$ are Borel. Note also that the Boolean operations are Borel over the space $\mathcal{K}(2^{\mathbb{N}})$.
It remains to verify that the assignments $\Gamma\mapsto K_{\Gamma}$ and $\Gamma\mapsto L_{\Gamma}$ are Borel. As in [CG01], we from a canonical enumeration of the clopen subsets of $2^{\mathbb{N}}$, denoted $\{\mathcal{O}^{\prime}_{i}\}_{i\in\mathbb{N}}$. To disjointify $K_{\Gamma}$ and $L_{\Gamma}$, we modify the enumeration to $\{\mathcal{O}_{i}\}_{i\in\mathbb{N}}$, where $\mathcal{O}_{2i}=\mathcal{O}_{2i+1}=\mathcal{O}^{\prime}_{i}$.
We first enumerate the clopen subsets of $E(\Gamma)$ and $E_{\ell}(\Gamma)$ respectively as $\{\mathcal{O}_{2i}\cap E(\Gamma)\}_{i\in\mathbb{N}}$ and $\{\mathcal{O}_{2i+1}\cap E_{\ell}(\Gamma)\}_{i\in\mathbb{N}}$, possibly with repetitions. Now let $\rho$ be the enumeration of the sequence $\{\mathcal{O}_{2i}\cap E(\Gamma)\}_{i\in\mathbb{N}}$ without repetitions (but preserving the order), and let $\rho_{\ell}$ be the enumeration of the sequence $\{\mathcal{O}_{2i+1}\cap E_{\ell}(\Gamma)\}_{i\in\mathbb{N}}$ without repetitions (but preserving the order). Then $\rho$ and $\rho_{\ell}$ are Borel functions that uniquely describe $K_{\Gamma}$ and $L_{\Gamma}$. ∎
We close this section by showing that $\operatorname{\thicksim_{\text{PHE}}}$ on $\mathcal{G}_{3}$ has a dense $G_{\delta}$ equivalence class, the class of one-ended graphs with infinite rank. This is the Loch Ness monster equivalence class, and the graphs in fig. 8 are examples.
Figure 8. Two representatives of the generic proper homotopy equivalence class of graphs in $\mathcal{G}_{3}$.
It is perhaps unsurprising that the generic graph is of infinite rank, since avoiding loops in the construction of a graph is quite restrictive. The key to showing that both finite rank and one-endedness are generic will be that rank and connectedness can be checked for inside balls of finite radius.
Proposition 3.3.
The equivalence class of one-ended graphs with infinite rank is dense $G_{\delta}$ in $\mathcal{G}_{3}$.
@@
First, we claim that the set of graphs in $\mathcal{G}_{3}$ with infinite rank is a dense $G_{\delta}$ set. Let $U_{n}$ be the graphs in $\mathcal{G}_{3}$ with rank at least $n$. Each set $U_{n}$ is open because for any $\Gamma\in U_{n}$, there is an $r\in\mathbb{N}$ such that $B_{r}(\Gamma)$ already has rank $n$. Then $\mathcal{U}_{(\Gamma,r)}\subseteq U_{n}$. To see that $U_{n}$ is dense, we show that for any $\Gamma\in\mathcal{G}_{3}$ and $r\in\mathbb{N}$, we can find $\Delta_{U}$ with rank at least $n$ and $B_{r}(\Delta_{U})\cong B_{r}({\Gamma})$. Such a graph $\Delta_{U}$ can be constructed from ${\Gamma}$ by subdividing $n$ edges of ${\Gamma}$ outside of $B_{r}({\Gamma})$ and attaching lollipop graphs to the new vertices. We finish the claim by observing that $\bigcap U_{n}$ is exactly the set of graphs of infinite rank.
Figure 9. For the black graph ${\Gamma}$, the ball of radius $2$ is drawn in blue. To find a one-ended graph in $\mathcal{U}_{({\Gamma},2)}$ we add the orange edge.
Because having infinite rank and one end are each generic properties and determine a unique PHE class, we see that the equivalence class of the Loch Ness monster graph is dense $G_{\delta}$. ∎
4. Extending the results to other spaces
We begin with a version of 3.1 for larger values of $k$.
Lemma 4.1.
For every $k>3$ and any closed subset $C$ of the Cantor space, $C$ is realized as the end space accumulated by loops of some $\Gamma_{C}^{k}\in\mathcal{G}_{k}$. Moreover, $\Gamma_{C}^{k}$ can be taken to have all of its ends accumulated by loops.
@@ 4. Extending the results to other spaces
We begin as in 3.1 with $C\subset 2^{\mathbb{N}}$, not a singleton, and the graph ${\Gamma}_{C}^{*}$. For each value of $k$ we describe how to modify ${\Gamma}_{C}^{*}$ to get a graph in $\mathcal{G}_{k}$ with $E({\Gamma}_{C}^{k})=E_{\ell}({\Gamma}_{C}^{k})=C$.
For $k$ odd, perform the following modifications to ${\Gamma}_{C}$:
  • for each degree $3$ vertex add $\frac{k-3}{2}$ self-loops,
  • for each degree $2$ vertex add a new vertex, connect them by $k-2$ edges, and add a self-loop to the new vertex.
This process adds loops to each vertex, so ultimately every end is accumulated by loops. See fig. 10 for an example where $k=7$.
Figure 10. Modifying the graph ${\Gamma}_{C}$ so that it is in $\mathcal{G}_{7}$.
For $k$ even, $k\neq 4$, we first double every edge of $\Gamma_{C}$ so that every vertex has degree $4$ or $6$. Then
  • for each degree $4$ vertex add $\frac{k-4}{2}$ self-loops,
  • for each degree $6$ vertex add $\frac{k-6}{2}$ self-loops.
Finally, the case where $k=4$ requires a bit more subtlety. First double every edge of ${\Gamma}_{C}$ and then perform a local PHE at every degree $6$ vertex to split it into two degree $4$ vertices, as in fig. 12. Because these are disjoint, compactly supported, homotopy equivalences, they do not change the PHE type of the graph, and in particular do not change the end space. Note that after doubling edges, every end of the graph was accumulated by loops.
Figure 12. A local PHE is done at each degree $6$ vertex to split it into two degree $4$ vertices. The choice of grouping of the $6$ edges into groups of $3$ does not affect the PHE type of the graph.
When $C$ is a singleton, we take ${\Gamma}_{C}^{*}$ to be a ray graph, so it has one vertex of degree $1$, call it $v$. To construct ${\Gamma}_{C}^{k}$, follow the instructions above with the following modifications for the degree one vertex.
  • For $k$ odd, add $\frac{k-1}{2}$ self-loops to $v$.
  • For $k$ even, after doubling edges $v$ will have degree $2$, so add $\frac{k-2}{2}$ self-loops.
The only place the $3$-regularity was used in the proof of 3.2 was 3.1. So, with 4.1 in hand, we have the complexity result for each space $\mathcal{G}_{k}$. Because $\mathcal{G}_{k}\subset\mathcal{G}_{\leq k}\subset\mathcal{G}_{\infty}$, we get the complexity result for each space $\mathcal{G}_{\star}$.
Corollary 4.2 (Intro 1.1 expanded).
Proper homotopy equivalence on $\mathcal{G}_{k}$, $\mathcal{G}_{\leq k}$, or $\mathcal{G}_{<\infty}$ is bireducible with $C_{\infty}$. That is, the equivalence relation $\operatorname{\thicksim_{\text{PHE}}}$ is Borel complete.
Next we show that the Loch Ness monster graph is still generic in the other spaces of infinite graphs.
Proposition 4.3 (Intro 1.2 expanded).
The equivalence class of one-ended graphs with infinite rank is dense $G_{\delta}$ in $\mathcal{G}_{k}$, $\mathcal{G}_{\leq k}$, and $\mathcal{G}_{<\infty}$.

### v2

@@ 4.1. Surfaces
By surface we mean a second countable, Hausdorff, orientable, 2-manifold without boundary. Two compact surfaces are homeomorphic if they have the same number of genus. The classification of non-compact surfaces predates, but is in direct analogy with, the classification of infinite graphs up to proper homotopy equivalence. It is due independently to Kerékjártó [Kerekjarto] and Richards [Richards].
@@
In this statement, $g(\Sigma)\in\mathbb{N}\cup\{0,\infty\}$ is the number of genus, $E(\Sigma)$ is the end space, and $E_{g}(\Sigma)$ is the subspace of ends accumulated by genus. The endspace of a surface can be defined mutatis mutandis with 2.1, replacing $\Gamma$} with $\Sigma$. To define the ends accumulated by genus, replace rank with genus in the definition of ends accumulated by loops.
4.2. Spaces of Surfaces
Next we want to realize sets of noncompact surfaces as Polish spaces; this includes infinite-type surfaces as well as finite-type surfaces with punctures. First, we define a pair of pants to be any compact, orientable, genus zero surface with at least one boundary component. If the pair of pants $P$ has $d+1$ boundary components we call them specifically a $d$-legged pair of pants, so that a $3$-holed sphere has two legs (the third boundary component being the waist). We allow $1$-legged pants, which are topologically annuli, and 0-legged pants, which are topologically disks. A pants decomposition of a surface $S$ is a collection of subsurfaces $\mathcal{P}=\{P_{j}\}_{j\in I}$ which are each themselves pants, cover the surface $S$, and overlap exactly on their boundary components. Denote the boundary components of $P_{j}$ by $\{\partial_{1}P_{j},\dots,\partial_{d+1}P_{j}\}$. That is, $S$ can be constructed by gluing the pants together along their boundary components by orientation preserving homeomorphisms, and we can write
@@ 4.2. Spaces of Surfaces
For each space of graphs $\mathcal{G}_{\star}$ discussed in Section 2.3 we will define an analogous space of surfaces, $\mathcal{S}_{\star}$. First, let $\hat{\mathcal{S}_{k}}$ be the set of surfaces with a (k1)(k-1)-legged pants decomposition. Similarly, let $\hat{S}_{\leq k}$ denote the set of surfaces with a pants decomposition in which all pants have at most k1k-1 legs, and $\hat{S}_{<\infty}$ denote the set of surfaces with any pants decomposition. For a surface with any type of pants decomposition $(S,\mathcal{P})$ we associate a graph in the following way. For each pair of paints $P_{j}$, associate a vertex $v_{j}$. Whenever a boundary component of $P_{j}$ is glued to a boundary component of $P_{k}$, associate an edge $(v_{j},v_{k})$, note that $j$ and $k$ may be equal. In this way, each (d1)(d-1)-legged pants corresponds to a vertex of degree $d$. If we consider surfaces with a basepoint interior to the pants decomposition, then we can associate a based graph. This association defines a function $\Phi_{\star}:\hat{\mathcal{S}}_{\star}\rightarrow\mathcal{G}_{\star}$ for each choice of $\star$. See fig. 10 for an example where the surface is compact.
[…]
@@
Each map $\Phi_{\star}:\hat{\mathcal{S}_{\star}}\rightarrow\mathcal{G}_{\star}$ is surjective, so we can then endow $\hat{\mathcal{S}_{\star}}$ with the initial topology. Recall that the initial topology is the coarsest topology which makes $\Phi_{\star}$ continuous and has the sets
$\mathcal{B}=\{\Phi_{\star}^{-1}(\mathcal{U})\mid\mathcal{U}\text{ is open in }\mathcal{G}_{\star}\}$
@@
as a subbase. Note that this topology will not be Hausdorff. For example, the same surface with two pants decompositions in which pants curves are isotopic, but not equal, are indistinguishable by open sets in the topology. So, we quotient by the equivalence relation induced by $\Phi_{\star}$, that is $(S,\mathcal{P},b)\sim(S^{\prime},\mathcal{P}^{\prime},b^{\prime})$ if $\Phi_{\star}(S,\mathcal{P},b)=\Phi_{\star}(S^{\prime},\mathcal{P}^{\prime},b^{\prime})$. Call these quotient spaces (with the quotient topologies) $\mathcal{S}_{\star}$, and note that they are Hausdorff and homeomorphic to the corresponding space $\mathcal{G}_{\star}$. For detailed explanation, see Corollary 22.3 of Munkres’ Topology, [munkres2000topology].
@@
Points in $\mathcal{S}_{\star}$ can be thought of as surfaces, up to homeomorphism, with based pants decompositions, up to homotopy. A puncture on a surface will map to a ray in the associated graph. So, similar to the discussion of which graphs live in which spaces $\mathcal{G}_{\star}$, there are no surfaces with isolated punctures in $\mathcal{S}_{k}$, and all homeomorphism types of noncompact surfaces are realized in $\mathcal{S}_{\leq k}$ for $k\geq 3$.
Let $\Psi_{\star}$ denote the induced homeomorphism from $\mathcal{S}_{\star}\rightarrow\mathcal{G}_{\star}$. By construction $\Psi_{\star}$ maps a genus to a loop, and an end to an end. As such, $\Psi_{\star}$ and it’s inverse are Borel reductions between proper homotopy equivalence and homeomorphism. From the results on $\mathcal{G}_{\star}$ we have the following.
Corollary 4.2 (Intro 1.4 expanded).
The equivalence relation of homeomorphism on each space $\mathcal{S}_{\star}$ is bireducible with $C_{\infty}$. In each space $\mathcal{S}_{\star}$ the equivalence class of surfaces with exactly one end which is accumulated by genus is comeager.
Appendix A The end space map is Borel
Here we carefully prove that the construction in 3.4 is Borel. We first prove that the map taking a finitely branching tree on $\mathbb{N}$ to its set of infinite branches is Borel. This is well known (see, e.g., [Kechris:classical_DST, Chapter 4]), but we include the proofs here for completeness. We remark that this map is not continuous (see, e.g., [Kechris:classical_DST, Exercise 4.32]).
Proposition A.1.
The space $\mathcal{T}_{f}$ of finitely branching trees on $\mathbb{N}$ is a Borel subset of $2^{\mathbb{N}^{<\mathbb{N}}}$.
@@ Appendix A The end space map is Borel
$\tau\in 2^{\mathbb{N}^{<\mathbb{N}}}$ is a finitely branching tree if and only if
  1. (1)
    $\tau(\emptyset)=1$.
  2. (2)
    For each $x\in\mathbb{N}^{<\mathbb{N}}$ and $i\in\mathbb{N}$ with $\tau(x^{\smallfrown}i)=1$, $\tau(x)=1$.
  3. (3)
    For each $x\in\mathbb{N}^{<\mathbb{N}}$, there is $N\in\mathbb{N}$ such that for all $i>N$, $\tau(x^{\smallfrown}i)=0$.
Condition (1) is open, (2) is $G_{\delta}$, and (3) is $\Pi^{0}_{3}$. ∎
Proposition A.2.
For any $x\in\mathbb{N}^{<\mathbb{N}}$ and corresponding cylinder set $[x]$, the set
$\mathcal{T}_{x}:=\{\tau\in\mathcal{T}_{f}:\;E(\tau)\cap[x]\neq\emptyset\}$
is Borel (in fact $G_{\delta}$) in $\mathcal{T}_{f}$.
@@
By König’s lemma, since each $\tau\in\mathcal{T}_{f}$ is finitely branching, $\tau$ has an infinite branch starting from $x$ if and only if the subtree below $x$ is infinite. Hence, $\mathcal{T}_{x}=\bigcap_{n\in\mathbb{N}}\bigcup_{y\supseteq x,|y|=n}\tau(y)=1$. ∎
Proposition A.3.
The end space map $E:\mathcal{T}_{f}\to\mathcal{K}(\mathbb{N}^{\mathbb{N}})$ that maps $\tau$ to its set of infinite branches is Borel.
@@
It is enough to show that for any open set $U\subseteq\mathbb{N}^{\mathbb{N}}$, the sets
$\mathcal{V}_{U}:=\{\tau\in\mathcal{T}_{f}:\;E(\tau)\cap U\neq\emptyset\}\text{ and }\mathcal{W}_{U}:=\{\tau\in\mathcal{T}_{f}:\;E(\tau)\subseteq U\}$
are Borel. Write $U$ as a countable union of cylinder sets, $U=\bigcup_{n\in\mathbb{N}}[x_{i}]$, where each $x_{i}\in\mathbb{N}^{<\mathbb{N}}$.
By A.2, $\mathcal{V}_{U}$ is $\Sigma_{2}^{0}$. Since each $E(\tau)$ is compact,
$\mathcal{W}_{U}=\bigcup_{n\in\mathbb{N}}\{\tau\in\mathcal{T}_{f}:\;E(\tau)\subseteq\bigcup_{i<n}[x_{i}]\}.$
Again by A.2, since $\bigcup_{i<n}[x_{i}]$ is clopen, each $\{\tau\in\mathcal{T}_{f}:\;E(\tau)\subseteq\bigcup_{i<n}[x_{i}]\}$ is $\Pi^{0}_{2}$, so $\mathcal{W}_{U}$ is $\Sigma^{0}_{3}$. Therefore, $E$ is a Borel map. ∎
We now show that the map taking a graph in $\mathcal{MG}_{\star}$ to its corresponding finitely branching tree on $\mathbb{N}$ (resp., its tree with rank $k$) from 3.4 is Borel.
Observation A.4.
For any subset $I\subseteq\mathbb{N}$, and any $i,n,m\in\mathbb{N}$, the following subsets of $\mathcal{MG}_{\star}$ are Borel. Below, by the ordering of the connected components, we mean the ordering induced by the usual order on $\mathbb{N}$. That is, we order the components by the labels of their smallest vertices.
  1. (1)
    $\mathcal{R}_{I}:=\{\Gamma\in\mathcal{MG}_{\star}:{\Gamma}[I]\text{ has positive rank}\}$
  2. (2)
    $\mathcal{P}_{I,i,j}:=\{\Gamma\in\mathcal{MG}_{\star}:i\text{ is connected to }j\text{ in }\Gamma[\mathbb{N}\setminus I]\}$
  3. (3)
    $\mathcal{C}_{I,n}:=\{\Gamma\in\mathcal{MG}_{\star}:\Gamma[\mathbb{N}\setminus I]\text{ has at least $n$ connected components}\}$
  4. (4)
[…]
@@
  1. (1)
    $\mathcal{R}_{I}$ is open, since this condition is witnessed on finitely many vertices.
  2. (2)
    $\mathcal{P}_{I,i,j}$ is open, since this condition is witnessed on finitely many vertices.
  3. (3)
    $\mathcal{C}_{I,2}$ is $F_{\sigma}$ for any $I$, since it is the countable union of complements of the $\mathcal{P}_{I,i,j}$. Similarly, each $\mathcal{C}_{I,n}$ is a countable Boolean combination of $\mathcal{P}_{I,i,j}$s.
  4. (4)
    $\Gamma\in\mathcal{C}_{I,n,m}$ if and only if
    • $\Gamma\in\mathcal{C}_{I,n}$, and
    • there are $m_{1},...,m_{n-1}<m$ such that the $m_{1},...,m_{n-1},m=m_{n}$ are pairwise disconnected outside of $I$ (that is, $\Gamma\in\bigcap_{i\neq j\leq n}\mathcal{P}_{I,m_{i},m_{j}}^{c}$), and
    • for all $j<m$, $j$ is connected to one of the $m_{i}$ outside of $I$ (that is, $\Gamma\in\bigcup_{i<n}\mathcal{P}_{i,j,m_{i}}$). ∎
Lemma A.5.
For each $x\in\mathbb{N}^{<\mathbb{N}}$, the set $\mathcal{H}_{x}:=\{{\Gamma}\in\mathcal{MG}_{\star}:\;x\in\tau_{\Gamma}\}$ is Borel. Hence, the map $\Gamma\mapsto\tau_{\Gamma}$, $\mathcal{MG}_{\star}\to{2}^{\mathbb{N}^{<\mathbb{N}}}$ constructed in 3.4 is Borel. Similarly, the map $\Gamma\mapsto\tau_{\Gamma}^{\ell}$ is also Borel.